Prenatal vitamins and Autism. Researchers studying the link between prenatal vitamin use and autism surveyed the mothers of a random sample of children aged 24 - 60 months with autism and conducted another separate random sample for children with typical development. The table below shows the number of mothers in each group who did and did not use prenatal vitamins during the three months before pregnancy.

Prenatal Vitamin Autism Typical Development Total
No vitamin 116 68 184
Vitamin 147 166 313
Total 263 234 497

Conduct the hypothesis test for the stated null and alternative. Use αα = K and round numeric answers to three decimal places.

H0H0: The variables use of prenatal vitamins during the three months before pregnancy and autism are independent.

HAHA: The variables use of prenatal vitamins during the three months before pregnancy and autism are not independent.

1. Enter the expected values for the hypothesis test in the table below, rounded to three decimal places.

Prenatal Vitamin Autism Typical Development
No vitamin
Vitamin

2. Calculate the test statistic for this hypothesis test.  X2X2 =

3. Calculate the degrees of freedom for this hypothesis test.

4. Calculate the p-value for this hypothesis test. p-value =  

5. What is your conclusion using αα = 0.01?

A. Do not reject H0H0
B. Reject H0H0
Answer #1

1)

Expected Ei=Σrow*Σcolumn/Σtotal Autism Typical development Total
No vitamin 97.368 86.632 184
Vitamin 165.632 147.368 313
Total 263 234 497

2)

chi square    χ2 =(Oi-Ei)2/Ei Autism Typical development Total
No vitamin 3.570 4.010 7.580
Vitamin 2.100 2.360 4.460
Total 5.670 6.370 12.040

  test statistic for this hypothesis test=12.040

3)

degree of freedom(df) =(rows-1)*(columns-1)= 1

4)

p-value for this hypothesis test=0.0005

5)B. Reject H0